# A Hierarchical Approach for Assessing the Vulnerability of Tree-Based Classification Models to Membership Inference Attack

### v1

@@ Abstract
More specifically, it is shown that the rank order of disclosure risk for different hyperparameter combinations remains consistent across datasets, enabling the development of simple, human-interpretable rules for identifying relatively high-risk models before training. While this ante-hoc analysis cannot determine absolute safety since this also depends on the specific dataset, it allows the elimination of unnecessarily risky configurations during hyperparameter tuning. Additionally, computationally inexpensive structural metrics serve as indicators of MIA vulnerability, providing a second filtering stage to identify risky models after training but before conducting expensive attacks. Empirical results show that hyperparameter-based risk prediction rules can achieve high accuracy in predicting the most at risk combinations of hyperparameters across different tree-based model types, while requiring no model training. Moreover, target model accuracy is not seen to correlate with privacy risk, suggesting opportunities to optimise model configurations for both performance and privacy.
@@ 1 Introduction
An increasing body of work has shown that machine learning (ML) models may expose confidential properties of the data on which they are trained [14]. This has resulted in a wide range of proposed attack methods with varying assumptions that exploit the model structure and/or behaviour to infer sensitive information, e.g., to perform model extraction, attribute inference, and membership inference [30, 19].
One of the most fundamental attacks is to identify whether an individual record was part of the training data, known as a membership inference attack (MIA). Recognising different sources of uncertainty in the model training process, the MIA can be framed as the question: “what are the probabilities that a model $M$ comes from the population $\mathcal{M}_{\text{in}}$ (or $\mathcal{M}_{\text{out}}$) of models trained with (or without) a particular record $X$?” For this to take a computational form requires a mapping from some set of metrics describing the structure and/or behaviour of models, and records, onto a probability distribution. Given the current state of knowledge, a key issue is how to identify one or more suitable metrics as proxies for underlying risk causes.
A common early approach uses the difference in confidence scores between predictions for samples that were in the training set and those that were not [33]. This approach fails to take into account that (hopefully) models will correctly identify regions where high probability records have the same label—whether they are from training set or not. The conceptual and methodological shortfalls of this approach have been well documented [28].
More recent attacks implicitly focus on decision boundaries by considering the record-wise difficulty of making correct predictions for the original task. For example, the Likelihood Ratio Attack (LiRA) uses as proxy the prediction error (loss) for a record $X$ from the original target model $M$, comparing it to the distributions of loss values from sampled subsets from $\mathcal{M}_{\text{in}}$ and $\mathcal{M}_{\text{out}}$ [7]. Similarly, the Quantile Regression [4] and Robust Membership Inference Attack (RMIA) [41] condition the confidence on that for other records (and models).
A number of other risk factors/proxy measures have been established. Overfitting (quantified by generalisation error) is well-known to be a significant cause of information leakage [39, 40]. Increasing model complexity therefore may pose a trade-off between accuracy and disclosure risk. For example, increasing the number of neural network parameters has been found to generally improve generalisation performance at the expense of lower privacy [37].
Consequently, a number of defences have been proposed [15, 17]. The most common of these is to incorporate information perturbation within the learning algorithm, such as adding noise through differential privacy (DP) [10]. However, the use of DP and knowledge distillation techniques may only mitigate some inference attacks [22]. Moreover, it has recently been highlighted that in practice DP implementations suffer from a number of problems and standard anti-overfitting techniques can often achieve better results [5]. For example, standard regularisation techniques have been shown to help mitigate the risk posed by an increasing number of model parameters [36].
The General Data Protection Regulation (GDPR) (and other laws, according to the data and jurisdiction) mean that before a ML model trained on confidential data can be released into the public domain, its vulnerability to attack should be assessed. However, some of the currently most effective attacks require considerable computational effort. For example, running a LiRA attack typically involves sampling (and hence training) 50–100 shadow models from $\mathcal{M}_{\text{out}}$. Therefore, from the perspective of people training models, and those tasked with that risk assessment, there is significant value in being able to rapidly identify vulnerable models.
@@
    This suggests that, like the conditioned loss metric at the heart of LiRA and other state-of-the-art attacks, they are proxies for some as yet unquantified latent property of models.
@@
These contributions allow us to avoid computationally costly MIA attacks and provide actionable insights into how to reduce disclosure risk.
@@ 2 Background
Trusted research environments (TREs) [^1] Also known as safe havens, secure data environments, and secure research services. offer a means for researchers to analyse confidential datasets and publish their findings. The globally recognised Five Safes approach to privacy preservation and regulatory compliance [12] requires TRE staff to independently check the disclosure risk posed by any proposed output before release is approved. This process of SDC is well understood for outputs of traditional statistical analyses. However, while the use of ML methods has rapidly expanded, a recent study of TREs by Kavianpour et al. [21] in the UK revealed a shortfall of SDC understanding and tools to support the use of ML. In particular, while researchers can already use ML methods within TREs, the inability to export those trained models precludes reproducibility of results, a critical tenet of scientific practice [3, 2]. After Fredrikson et al. [11] demonstrated that ML models can be vulnerable to attacks inferring aspects of training data, there has been a rapid growth in proposed attacks and counter-measures as surveyed by Hu et al. [14]. Regulatory authorities such as the UK Information Commissioner’s Office [16] recognise this risk, but offer general advice rather than practical guidance and tools.
@@
A recent report by DARE UK [9] highlighted the need to “Where possible, automate the review of outputs to support and focus the use of skilled personnel on the areas of most significant risk”. Moreover, Jefferson et al. [18] have recently presented a detailed discussion of the implications of ML model release and proposed a set of guidelines for TREs.
@@ 2.1.1 TREs and SDC
Nowadays, national statistics institutes (and other holders of confidential information such as health authorities) routinely conduct SDC of all outputs they produce. They also offer access to researchers via TREs. TREs almost universally employ the Five Safes [^2] Safe people, safe projects, safe settings, safe data, safe outputs. framework to provide multiple layers of assurance to data owners, within which the final stage of checking for ‘safe outputs’ is a vital part.
@@ 2.1.2 Implications for ML Privacy Research
Thus for traditional forms of analyses, there is decades of well understood theory and practice surrounding the SDC of outputs such as (multi-dimensional) tables, regression models, survival statistics and a range of other outputs [6, 13] which have recently been consolidated into a coherent ‘statbarns’ framework [31]. From these we can identify three key concepts relevant to ML models:
@@
  • Class disclosure risk. It should not be possible to make inference about all members of a subgroup via the publication of query results containing zeros, e.g., “in our fictitious survey no ML researchers were aware of the Secure Data Access Professionals manual [13] [^3] In some cases it may be argued that zeros may be structural (people born without certain organs are unlikely to develop cancers of those organs) or may be ‘evidential’ and scientifically important: “no people with gene XYZ had disease A”. Hence the practice of applying principles-based SDC rather than strictly applying rules..
@@ 3 Methodology
3.1 Attacks and Performance Metrics
LiRA [7] is used here to perform membership inference and assess the disclosure risk presented by a range of fitted tree-based classification models. LiRA is a fine-grained black-box evaluation metric-based MIA [24] and (especially when using sufficient shadow models) is representative of the current state-of-the-art. Moreover, it places an emphasis on an attacker scenario that corresponds to TRE concerns: “If a membership inference attack can reliably violate the privacy of even just a few users in a sensitive dataset, it has succeeded”.
Building on this line of thinking, we note that attackers will almost always have access to some records they know are non-members of the training set—either because they have artificially created them, or perhaps because they may be their own records. We posit that armed with that information, and aware (from the literature) that in most cases the MIA cannot make a confident prediction, attackers will prefer a model that: (i) does not make high confidence member predictions for records they know are non-members; (ii) is correct when it predicts non-member with high confidence.
From these premises we derive a frequency difference metric $\text{FDIF}_{z}$, calculated as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
The attack uses 100 shadow models to estimate the loss distributions. For each target model hyperparameter combination, 10 random training and test data splits are performed and a LiRA attack executed. Since the ML privacy community has yet to reach a consensus on a single disclosure risk metric, here a range of metrics are computed. The results are therefore averaged over the 10 runs to produce a set of risk metrics for a given target model hyperparameter combination on a dataset.
In addition to several MIA success metrics, several target model metrics are also examined for correlations. These are described in Table LABEL:table:metrics.
3.2 Statistical Significance of Attack Vulnerabilities
@@ 3.2.1 Preliminaries
MIAs may be created using synthetic datasets which approximate the underlying distribution from which $\mathcal{D}$ is drawn. However, as reported by Carlini et al. [7] and Shokri et al. [33], and especially Ye et al. [38], the success of those attacks increases as sources of uncertainty in that approximation are reduced or removed.
Thus, in recognition of the pace at which new MIAs are being invented, in our worst-case scenario, we remove those uncertainties to estimate an upper bound on the risk. Hence we explicitly use the sets of outputs (either directly as probabilities or converted to losses) as the basis for conducting different MIAs.
  • Note that without loss of generality we assume that an attack creates an MIA regression model $A:M(\mathcal{D})\rightarrow\mathcal{R}$. Given a target model output $x^{\prime}=M(x),x\in\mathcal{D}$, we condition the outputs so that $A(x^{\prime})$: represents the estimated probability that $x\in\mathcal{D}_{\text{train}}$.
@@ 3.2.2 From True Positive Rates to Attack Vulnerability
However, from the perspective of an SDC output checker tasked with making a decision on whether to approve the release of the trained model $M$, what would be more helpful is to quantify the likelihood of observing different numerical MIA classifier results (TPR@$z$) by chance. We assess this via a null hypothesis:
@@
Thus after some rearrangement we assert that with $(1-\alpha)\times 100$% confidence, a model $M$ is LiRA vulnerable at a FPR of $z$ if:
@@
3.3 Target Models

### v2

@@ Abstract
More specifically, it is shown that the rank order of disclosure risk for different hyperparameter combinations remains consistent across a range of datasets, enabling the development of simple, human-interpretable rules for identifying relatively high-risk models before training. While this ante-hoc analysis cannot determine absolute safety since this also depends on the specific dataset, it allows the elimination of unnecessarily risky configurations during hyperparameter tuning. Additionally, computationally inexpensive structural metrics serve as indicators of MIA vulnerability, providing a second filtering stage to identify risky models after training but before conducting expensive attacks. Empirical results show that hyperparameter-based risk prediction rules can achieve high accuracy in predicting the most at risk combinations of hyperparameters across different tree-based model types, while requiring no model training. Moreover, target model accuracy is not seen to correlate with privacy risk, suggesting opportunities to optimise model configurations for both performance and privacy.
@@ 1 Introduction
An increasing body of work has shown that machine learning (ML) models may expose confidential properties of the data on which they are trained [41]. This has resulted in a wide range of proposed attack methods with varying assumptions that exploit the model structure and/or behaviour to infer sensitive information. For example, to perform data reconstruction, model extraction, attribute inference, and membership inference [14, 20].
The General Data Protection Regulation (GDPR) (and other laws, according to the data and jurisdiction) mean that before an ML model trained on confidential data can be released into the public domain, its vulnerability to attack should be assessed [51]. One of the most fundamental techniques to assess ML privacy disclosure is the membership inference attack (MIA) [45], which attempts to identify whether an individual record was part of the original training data. Since many of the other attacks have been shown to be more difficult to perform [57], vulnerability to MIA can be seen as a strong indicator of potential privacy risks, and its presence suggests that further, more sophisticated attacks might also be feasible, even if more challenging to execute. Therefore, a model’s robustness against MIA is a crucial first step in demonstrating its privacy-preserving capabilities before public deployment.
However, some of the currently most effective MIAs require considerable computational effort. For example, running a Likelihood Ratio Attack (LiRA) [5] typically involves sampling (and hence training) 50–100 shadow models. Therefore, from the perspective of people training models, and those tasked with that risk assessment, there is significant value in being able to rapidly identify vulnerable models or reduce the number that require computationally expensive evaluation.
@@
    This suggests that the structural metrics are proxies for some as yet unquantified latent property of the models, similar to the conditioned loss metric at the heart of LiRA and other state-of-the-art attacks.
@@
These contributions allow us to avoid computationally costly MIA attacks and provide actionable insights into how to reduce disclosure risk. The remainder of this paper is organised as follows. We begin in Section 2 by discussing MIAs, traditional SDC within TREs, and the implications for ML privacy. Section 3 then details the methodology used to investigate the privacy risks of hyperparameters and structural metrics in tree-based ML models. We present our empirical results in Section 4, and finally, Section 5 summarises and discusses our findings.
@@ 2 Background
2.1 Membership Inference Attacks
Recognising different sources of uncertainty in the model training process, the MIA can be framed as the question: what are the probabilities that a model $M$ comes from the population $\mathcal{M}_{\text{in}}$ (or $\mathcal{M}_{\text{out}}$) of models trained with (or without) a particular record $X$? For this to take a computational form requires a mapping from some set of metrics describing the structure and/or behaviour of models, and records, onto a probability distribution. Given the current state of knowledge, a key issue is how to identify one or more suitable metrics as proxies for underlying risk causes.
A common early approach to MIA exploits the difference in confidence scores between predictions for samples that were in the training set and those that were not [45]. However, this approach fails to take into account that (hopefully) models will correctly identify regions where high probability records have the same label—whether they are from training set or not. The conceptual and methodological shortfalls of this approach have been well documented [39].
More recent attacks implicitly focus on decision boundaries by considering the record-wise difficulty of making correct predictions for the original task. For example, LiRA uses as proxy the prediction error (loss) for a record $X$ from the original target model $M$, comparing it to the distributions of loss values from sampled subsets from $\mathcal{M}_{\text{in}}$ and $\mathcal{M}_{\text{out}}$ [5]. Similarly, the quantile regression [3] and robust MIA (RMIA) [56] condition the confidence on that for other records (and models). Relative measures of MIA vulnerability have also been suggested, computing risk based on the difference in prediction loss for member samples relative to a predefined reference model [2]. Recent work has also sought to improve performance with target model knowledge distillation [11].
MIAs typically assess whether a given data point exactly matches a training record, which potentially neglects similar or overlapping memorised data. For example, an image from a different angle may still be recognisable and evoke an exploitable behaviour from the model even though it would be incorrectly assessed as a false positive with a binary exact matching metric. Consequently, Tao and Shokri [50] propose a broader assessment framework that evaluates whether the model was trained on any data within a specified range using domain-specific distance measures.
Typically, MIAs assume that the model’s confidence scores (or loss values) are exposed. However, even when this information is hidden or unavailable, label-only based attacks can still identify whether a record was in the original training data. This is achieved by perturbing records to assess model robustness with the assumption that records exhibiting high robustness belong to the training dataset [7]. Moreover, in some cases training membership can be accurately inferred with just a single query [32] by exploiting the fact that member samples exhibit more resistance to adversarial perturbations than non-member samples.
One of the most significant drawbacks for assessing ML privacy with MIAs is their computational cost, particularly when using shadow models, which can require training many models with the same architecture and algorithm as the original target model. This can mean that they may not be feasible for large models or datasets. Consequently, a growing body of work is exploring ways to reduce the computational burden, for example maintaining MIA accuracy with only a few shadow models [56]. Interestingly, a recent theoretical analysis of likelihood ratio-based MIAs on stochastic gradient descent [23] has shown how various hyperparameters influence vulnerability, without needing to train any shadow models.
A number of risk factors and proxy measures have been established, with overfitting (quantified by generalisation error) recognised as a significant cause of information leakage [54, 55]. Increasing model complexity often entails a trade-off between privacy and utility, as observed in neural networks where more parameters can enhance performance but compromise privacy [49]. While explainable ML models are often proposed to boost transparency and trust, recent studies caution that they may inadvertently increase vulnerability by revealing feature importance information, which can be exploited [24, 27]; perturbing these features typically results in a bigger loss in confidence for member samples.
Implementing robust ML practices to prevent overfitting and ensure generalisation is therefore an important first step in mitigating MIA vulnerabilities [9]. Additionally, a number of defences have been proposed [15, 18]. The most common of these is to incorporate information perturbation within the learning algorithm, such as through differential privacy (DP) [10]. However, the use of DP and knowledge distillation techniques may only mitigate some inference attacks [25]. Indeed it has been suggested that model distillation may not provide any effective privacy defence [16], and the level of DP necessary to protect a model from MIA may render its utility unusable [37].
Moreover, it has recently been highlighted that in practice DP implementations suffer from a number of problems and standard anti-overfitting techniques can often achieve better results [4]. For example, standard regularisation techniques have been shown to help mitigate the risk posed by an increasing number of model parameters [48].
As the use of ML continues to proliferate, it is becoming increasingly common to release multiple updated versions of a model. However, doing so may further expose confidential properties of the data, strengthening reconstruction attacks [44]. This has also been shown to be true for improving MIA success [17], which may be a particular risk for secure data facilities such as TREs that may egress many models trained on the same data, and is a similar concern to differencing attacks in traditional output SDC.
2.2 Trusted Research Environments
Trusted research environments (TREs) [^1] Also known as safe havens, secure data environments, and secure research services. offer a means for researchers to analyse confidential datasets and publish their findings. The globally recognised Five Safes [^2] Safe people, safe projects, safe settings, safe data, safe outputs. approach to privacy preservation and regulatory compliance [13] requires TRE staff to independently check the disclosure risk posed by any proposed output before release is approved.
This process of SDC is well understood for outputs of traditional statistical analyses. However, while the use of ML methods has rapidly expanded, a recent study of TREs by Kavianpour et al. [22] in the UK revealed a shortfall of SDC understanding and tools to support the use of ML. In particular, while researchers can already use ML methods within TREs, the inability to export those trained models precludes reproducibility of results, a critical tenet of scientific practice [1].
@@ 2.2 Trusted Research Environments
Regulatory authorities such as the UK Information Commissioner’s Office recognise the risks of ML models, but offer general advice rather than practical guidance and tools. Jefferson et al. [19] have recently presented a detailed discussion of the implications of ML model release and proposed a set of guidelines for TREs. A recent report by DARE UK [8] highlighted the need to “Where possible, automate the review of outputs to support and focus the use of skilled personnel on the areas of most significant risk”.
@@ 2.3.1 TREs and SDC
Nowadays, national statistics institutes (and other holders of confidential information such as health authorities) routinely conduct SDC of all outputs they produce. They also offer access to researchers via TREs. TREs almost universally employ the Five Safes framework to provide multiple layers of assurance to data owners, within which the final stage of checking for ‘safe outputs’ is a vital part.
@@ 2.3.2 Implications for ML Privacy Research
Thus for traditional forms of analyses, there is decades of well understood theory and practice surrounding the SDC of outputs such as (multi-dimensional) tables, regression models, survival statistics and a range of other outputs, which have recently been consolidated into a coherent ‘statbarns’ framework [12]. From these we can identify three key concepts relevant to ML models:
@@
  • Class disclosure risk. It should not be possible to make inference about all members of a subgroup via the publication of query results containing zeros, e.g., “in our fictitious survey no ML researchers were aware of the Secure Data Access Professionals manual” [^3] In some cases it may be argued that zeros may be structural (people born without certain organs are unlikely to develop cancers of those organs) or may be ‘evidential’ and scientifically important: “no people with gene XYZ had disease A”. Hence the practice of applying principles-based SDC rather than strictly applying rules..
@@ 3 Methodology
This section presents the methodology used to investigate MIA risks in tree-based classification models. We first state our hypotheses for disclosure risk based on model hyperparameters in Section 3.1 and structural model properties in Section 3.2. Section 3.3 provides further details of the statistical framework used for determining the significance of observed MIA success. We then describe our experimental setup in Section 3.4. This includes details of the specific MIA and metrics explored, including a novel metric $\text{FDIF}_{z}$, as well as target models and datasets used for evaluation.
3.1 Disclosure Risk and Model Hyperparameters
It is here hypothesised that while absolute disclosure risk is dataset-dependent, the relative risk ranking of different model hyperparameter combinations may generalise across datasets. However, since many different risk metrics have been proposed, it is first necessary to (i) select a single representative measure to analyse the relative risk; and (ii) establish that this metric is likely to generalise across different datasets. Subsequently, we test whether a simple, interpretable model can predict if a hyperparameter combination leads to relatively high or low risk as defined by this chosen metric.
That is, we hypothesise:
  • H1.1: Since MIA metrics measure different aspects of the same phenomenon (risk), hyperparameter rank orderings for one metric will be predictive of rank orderings for another. To test this hypothesis, for each target model algorithm, pair-wise rank correlation statistics are computed to compare the relative risk between metrics across a range of hyperparameter combinations for each dataset.
  • H1.2: This rank ordering will generalise across datasets. To test this hypothesis, we select a single representative MIA metric and compute Kendall $\tau$ rank correlation coefficients of hyperparameter risk to perform pair-wise comparisons for each dataset.
  • H1.3: We can exploit H1.1 and H1.2 to produce a risk prediction model for ‘unnecessary vulnerability’ that can be applied to different datasets without the need for sampling MIA scores on these new datasets. To test this hypothesis, we assess the performance of models trained to predict whether a given hyperparameter combination is relatively high or low risk as a binary classification task. These models are trained on a single dataset (mimic2-iaccd) that has been sampled for MIA risk, and the rules are then extracted and evaluated on the remaining (unseen) datasets to test generalisation performance.
    That is, for each target model type, a Decision Tree classifier is used to learn a simple human-interpretable ruleset that predicts whether a hyperparameter combination is relatively high or low risk. Risk is here defined as high if the previously chosen metric is in the top 20% of all hyperparameter combinations. While this is a somewhat arbitrary threshold, the aim is to prove that a simple set of rules can be identified that predict whether a hyperparameter combination poses unnecessarily high risk.
    In order to extract rules that are easy to interpret, the Decision Trees are restricted to a maximum depth of 5; using a simple grid search for the other hyperparameters. Stratified 10-fold cross-validation was applied to the single training dataset (mimic2-iaccd) and the best performance was found in each case to be a minimum number of samples per split of 2, a minimum number of samples per leaf of 2, the entropy criterion, and a balanced class weight.
3.2 Disclosure Risk and Structural Model Properties
To investigate the relationship of disclosure risk and the structural model properties described in Section 2.3.2, we take the following approach. For a range of algorithm-dataset-hyperparameter combinations sampled, we analyse the model produced and its predictions for the training set, and use these to capture the structural metrics.
We then join these with the various metrics captured from the MIAs and use this combined data to examine two complementary hypotheses:
  • H2.1: Structural risks are necessary conditions for model vulnerability. In other words, if a given structural metric is 0 then there will be no MIA risk. In the following analysis we insist on 100% support for hypothesis 2.1 because this is a statement about safety.
  • H2.2: The presence of structural risks are sufficient conditions for model vulnerability. In other words, if (the product of a combination of) structural metric(s) is 1 there will be non-zero MIA risk. We insist on high, but not 100% support for hypothesis 2.2 because this is a statement about risk.
3.3 Statistical Significance of MIA Vulnerabilities
@@ 3.3.1 Preliminaries
MIAs may be created using synthetic datasets which approximate the underlying distribution from which $\mathcal{D}$ is drawn. It is well-established that the success of MIAs often increases as uncertainties in approximating the data distribution are reduced [5, 45, 53].
Thus, in recognition of the pace at which new MIAs are being invented, in our scenario, we remove those uncertainties to estimate an upper bound on the risk. Hence we explicitly use the sets of outputs (either directly as probabilities or converted to losses) as the basis for conducting different MIAs.
For the purpose of evaluating the statistical significance of attack vulnerabilities, we consider a generic MIA regressor $A$ derived from the target model $M$’s outputs. That is:
  • Without loss of generality, we assume that an attack creates an MIA regression model $A:M(\mathcal{D})\rightarrow\mathcal{R}$. Given a target model output $x^{\prime}=M(x),x\in\mathcal{D}$, we condition the outputs so that $A(x^{\prime})$: represents the estimated probability that $x\in\mathcal{D}_{\text{train}}$.
@@ 3.3.2 From True Positive Rates to Attack Vulnerability
However, from the perspective of an SDC output checker tasked with making a decision on whether to approve the release of the trained model $M$, what would be more helpful is to quantify the likelihood of observing different numerical MIA classifier results (TPR@$z$) by chance.
We assess this via a null hypothesis:
@@
Thus after some rearrangement we assert that with $(1-\alpha)\times 100$% confidence, a model $M$ is MIA vulnerable at a FPR of $z$ if:
@@
3.4 Experimental Setup
3.4.1 Membership Inference Attack
Here we use the LiRA [5] implementation from the open source SACRO-ML framework [47] to perform membership inference and assess the disclosure risk presented by a range of fitted tree-based classification models. LiRA is a fine-grained black-box evaluation metric-based MIA [30] and (especially when using sufficient shadow models) is representative of the current state-of-the-art. Furthermore, it places an emphasis on an attacker scenario that corresponds to TRE concerns: “If a membership inference attack can reliably violate the privacy of even just a few users in a sensitive dataset, it has succeeded”. For each target model hyperparameter combination, 10 random training and test data splits are performed to fit 10 different models, and a LiRA attack is executed for each. The attack uses 100 shadow models to estimate the loss distributions.
3.4.2 Disclosure Metrics
Since the ML privacy community has yet to reach a consensus on a single disclosure risk metric, a range of standard MIA success metrics are computed from the LiRA outputs; for example, TPR at low FPR, and area under the curve (AUC). These metrics are averaged over the 10 runs to produce a set of risk metrics for a given target model hyperparameter combination on a dataset.
Moreover, we note that attackers will almost always have access to some records they know are non-members of the training set—either because they have artificially created them, or perhaps because they may be their own records. We therefore posit that armed with that information, and aware (from the literature) that in most cases the MIA cannot make a confident prediction, attackers will prefer a model that: (i) does not make high confidence member predictions for records they know are non-members; (ii) is correct when it predicts non-member with high confidence.
From these premises we derive a novel frequency difference metric $\text{FDIF}_{z}$, calculated as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
In addition to these MIA-specific metrics, several target model metrics (e.g., generalisation error) are also examined for correlations. All metrics considered here are described in Table LABEL:table:metrics.
3.4.3 Target Models